What mass of Al(OH)3 is produced if 22.7g of NaOH is consumed?

1
Expert's answer
2022-04-28T05:49:05-0400

Solution:

The molar mass of NaOH is 40 g/mol

Therefore,

Moles of NOH = [22.7 g NaOH] × [1 mol NaOH / 40 g NaOH] = 0.5675 mol NaOH


Balanced chemical equation:

Al2(SO4)3 + 6NaOH → 3Na2SO4 + 2Al(OH)3

According to stoichiometry:

6 mol of NaOH produce 2 mol of Al(OH)3

Thus, 0.5675 mol of NaOH produce:

[0.5675 mol NaOH] × [2 mol Al(OH)3 / 6 mol NaOH] = 0.1892 mol Al(OH)3


The molar mass of Al(OH)3 is 78 g/mol

Therefore,

Mass of Al(OH)3 = [0.1892 mol Al(OH)3] × [78 g Al(OH)3 / 1 mol Al(OH)3] = 14.76 g Al(OH)3

Mass of Al(OH)3 = 14.76 g = 14.8 g


Answer: 14.8 grams of Al(OH)3 are produced



Short form solution:

Balanced chemical equation:

Al2(SO4)3 + 6NaOH → 3Na2SO4 + 2Al(OH)3

According to stoichiometry:

Mass of Al(OH)3 = [22.7 g NaOH] × [1 mol NaOH / 40 g NaOH] × [2 mol Al(OH)3 / 6 mol NaOH] × [78 g Al(OH)3 / 1 mol Al(OH)3] = 14.755 g Al(OH)3 = 14.8 g Al(OH)3